OpenStudy (juan1857):

Find each principle root. Express the result in the form a+bi with a and b rounded to the nearest tenth: (2+2i)^1/3

OpenStudy (loser66):

Where are you stuck?

OpenStudy (juan1857):

I only went up to \[(2\sqrt{2})^{1/3}(\cos45+360(0)+isin46+360(0))\]

OpenStudy (loser66):

Let z= 2+2i, then we need find third roots of it. \(|z|=2\sqrt2\)

OpenStudy (loser66):

and \(z=|z|e^{i\theta}\)

OpenStudy (loser66):

Let \(\omega\) be roots of z, and \(\omega =|\omega| e^{i\tau}\) \(\omega^3=|\omega|^3e^{3i\tau}\)

OpenStudy (loser66):

\(|\omega|^3= 2\sqrt 2\), hence \(|\omega| =\sqrt[6]8\)

OpenStudy (loser66):

\(e^{3i\tau}= e^{i\theta}\), then \(3\tau = \theta +2\pi n\), n =0,1,2 That gives us \(\tau = \theta /3+2\pi n/3\)

OpenStudy (loser66):

what is \(\theta\)? 

OpenStudy (loser66):



OpenStudy (loser66):

Hence \(\theta =\pi/4\)